# GeoTransformer: Enhancing Urban Forecasting with Geospatial Attention Mechanisms

### v1

@@ Introduction
We train various baseline models for comparison. As detailed in Table 1, models in Panel 1 utilize either direct numeric or imagery input. Specifically, Model 1 employs a linear regression configured with 12-dimensional sociodemographic data as input. Model 2 utilizes a pre-trained Visual Transformer (ViT) (Dosovitskiy et al. 2020a), which is subsequently fine-tuned on the prediction tasks with the last 4 encoder layers and the output layer unfrozen. Panel 2 presents models that utilize latent representations as inputs to perform predictions. Following DHM(Wang et al. 2024), model 3 employs a polynomial regression. Model 4 adopts a Graph Attention Network (GAT) (Zhang, Yu, and Liu 2019) model with each city region as a node connected to its neighbors, forming the graph structure. All hyperparameters for training the baseline models are optimized through a grid search to ensure peak performance on the test set.
@@
As shown in Table 1 panel 3, we train the GeoTransformer (model 5-7) using three spatial weighting methods. The number of decoder layers is set to 4, and the number of attention heads is set to 16. For each model variant and task, the L2 regularization decay weight was optimized, with a range of 0.0001 to 0.1, while the number of neighboring regions is set to 49, 81, or 169. All parameters are grid-searched to ensure the best performance. For both GDP and ride-share demand predictions, mean squared error (MSE) is employed as the loss function. All models are trained on an RTX 4090 GPU running Ubuntu 22.04.
@@
Table 2 summarizes the predictive performance evaluation of baselines and GeoTransformer models in both GDP and ride-share demand prediction tasks.
@@
As shown in Table 2, all three GeoTransformer models surpass every baseline model. Compared to Random Forest and ViT in panel 1, our best model (linear weighting) outperforms by 27% and 18% respectively in $R^{2}$. In Panel 2, the GAT model exhibits robust decoding capabilities for urban latent representations and delivers the best predictive accuracy among baseline models, while our GeoTransformer with linear weighting surpasses it by 14% in $R^{2}$. This demonstrates that our geospatial attention mechanism effectively models spatial dependencies without the need for predefined graph structures. In panel 3, among the three GeoTransformer models operating under identical parameters and training epochs, the linear weighting approach exhibits the best predictive performance.
@@
The results in Panel 1 of Table 2 indicate that ride-share demand is more strongly influenced by sociodemographic factors. Both polynomial regression and GAT models struggle to effectively decode latent representations for accurate predictions. However, our GeoTransformer model with IDW weighting outperforms the best baseline by 18% in $R^{2}$ on the test set. Additionally, among the GeoTransformer variants, it is observed that the IDW model not only achieves the highest performance but also exhibits the fastest convergence.
Figure 4: An example of the attention map between the query region and its neighboring value regions.
Qualitative Analysis.
To investigate the functionality of the geospatial attention mechanism, we visualize several query regions’ attention maps with their neighboring value regions. We present an example in Figure 4, which (a) includes satellite images of the northwest area of the Gresham community, with the query region marked by a red marker (Figure 4), and its attention map for GDP prediction. The query region focuses on its own community area, nearby amenities (the park in the lower left and the supermarket and church in the upper right), and critical road infrastructure (the freeway entrance in the upper left). These results validate the effectiveness of the geospatial attention mechanism in accurately capturing urban features relevant to GDP.
@@
To investigate the necessity and effectiveness of each design in GeoTransformer, we conduct ablation experiments for geospatial weighting, trainable key matrices, and the number of neighboring value regions.
Figure 5: The GDP prediction accuracy comparison on the test set between no weighting attention and geospatial weighted attention during the training process.
Geospatial Attention Module.
To demonstrate the effectiveness of the geospatial attention module, we conducted a controlled experiment by training an alternative version of the model devoid of the weighting attention mechanism. This baseline model maintained the same architecture and training parameters as the GeoTransformer models, featuring 16 heads, 81 neighbors, 4 layers, and an L2 regularization decay weight of 0.0001, trained over 50 epochs. As shown in Figure 5, the comparative analyses illustrate that all three geospatial attention variants significantly enhanced the GDP prediction accuracy.
Trainable Key Matrices.
Intuitively, the correlation between different regions in urban forecasting is not directly determined by the similarity between these regions. Therefore, we set the key matrix in the attention mechanism as trainable parameters to flexibly adjust the correlation between the value regions and the query region. To validate the effectiveness of this design, we train a GeoTransformer model by setting key matrices $K$ fixed and identical to value matrices $V$, while keeping all other settings constant. As shown in Table 3, linear weighting GeoTransformer with trainable $K$ achieves higher prediction accuracy over 5% in $R^{2}$. It is also observed that the fixed $K$ GeoTransformer converges faster during training, which is reasonable given its smaller number of trainable parameters. However, the flexibility of a trainable $K$ allows the model to ultimately achieve higher accuracy.
[…]
Fixed $K$ 0.241/0.410 0.364/0.464 0.936/0.893
Trainable $K$ 0.081/0.216 0.210/0.346 0.979/0.944
Table 3: Comparison with linear weighting GeoTransformer with fixed $K$ matrix or trainable $K$ matrix.
$k$-Nearest Neighbors.
The spatial extent of surrounding information required by the query region may vary for different urban forecasting tasks. Therefore, we search for the optimal number of neighboring regions $k$ for GDP and ride-share demand predictions. By centering on the query region and expanding outward in a square, $k$ is determined by the square of the side length. We linearly increase the side length to identify the optimal $k$ that maximizes model performance.
For GDP prediction, providing the model with a broader range of urban information initially leads to an improvement in accuracy. However, as $k$ increases further, the model may become overwhelmed by excessive information, resulting in a decline in accuracy (Figure 6).
A similar trend is observed in the ride-share demand prediction, where increasing $k$ initially enhances model accuracy. The model performs best when $k$ is set to 49 or 81. However, further increasing $k$ by incorporating additional neighboring information leads to a decline in accuracy (Figure 6).
Figure 6: Prediction accuracy comparison across $k$ nearest neighbors.
@@
We propose a GeoTransformer structure to extract information from high-dimensional urban latent representations. The novel geospatial attention mechanism allows GeoTransformer to capture extensive urban information and its spatial dependency. The experimental results demonstrate the capacity of our method in urban forecasting tasks. An ablation study is conducted to validate the effectiveness of the geospatial attention mechanism and the design of GeoTransformer.
Self-attention module in Transformer (Vaswani 2017) is primarily intended for sequence prediction tasks. In urban forecasting, spatiotemporal prediction tasks also resemble sequence prediction. Although this study does not activate the self-attention module, we retain it in the GeoTransformer as it holds the potential for adapting to spatiotemporal sequence prediction tasks. This merits further exploration in subsequent research.

### v2

@@ Introduction
We train various baseline models for comparison. As detailed in Table 1, models in Panel 1 utilize the deep hybrid model (DHM) (Wang et al. 2024) framework, directly inputting representations to regression models to predict. Specifically, Model 1 employs a linear regression, and Model 2 utilizes a polynomial regression. Panel 2 presents models that utilize representations as inputs for graph attention networks (GATs). Model 3 forms the graph structure by assigning each region as a node connected to its rook neighbors. Model 4 constructs a graph using the sparse matrix, where each region is a node connected to its top-$k$ relevant neighbors selected by the highest weights in our sparse matrix, with edges weighted accordingly. The number of top-$k$ connections in Model 4 is consistent with the retrieval size in GeoTransformer, while all other hyperparameters for training the baseline models are optimized through a grid search to ensure peak performance on the test set.
@@
As shown in Table 1 panel 3, we train the GeoTransformer (model 5) with the number of decoder layers set to 4, and the number of attention heads set to 16. The L2 regularization decay weight is optimized at 0.001, while the number of retrieved regions is set to 81. All parameters are grid-searched to ensure the best performance. For both GDP and ride-share demand predictions, mean squared error (MSE) is employed as the loss function. All models are trained on an RTX 4090 GPU running Ubuntu 22.04.
@@
Table 2 summarizes the predictive performance evaluation of baselines and the GeoTransformer model in both GDP and ride-share demand prediction tasks.
@@
As shown in Table 2, our GeoTransformer model surpasses every baseline model. For Deep Hybrid Models in panel 1, linear regression performs a little better than polynomial regression, and our model outperforms them by around 80% in $R^{2}$. In Panel 2, the GAT models exhibit similar capabilities for decoding region-based representations to predict GDP, with over 80% in $R^{2}$, while our GeoTransformer still surpasses them by 10% in $R^{2}$.
[…]
1 Random Regions + No Attention Weighting 3.24/3.84 1.54/1.65 0.061/0.007
2 Retrieved Module + No Attention Weighting 0.618/0.650 0.596/0.708 0.811/0.758
3 Retrieved Module + Spatial Proximity Weighting 0.245/0.418 0.364/0.478 0.906/0.881
4 Retrieved Module + Spatial Proximity & Information Entropy Weighting 0.079/0.216 0.207/0.343 0.964/0.933
Table 3: Ablation experimental results of each GeoTransformer component.
@@
The results in Panel 1 of Table 2 show that both Deep Hybrid Models achieve 60% and 33% in $R^{2}$ respectively, while GeoTransformer surpasses them over 33 % in $R^{2}$. In panel 2, the GAT model based on our sparse matrix achieves nearly 80% in $R^{2}$, around 22% higher than the one with grid structure. However, our GeoTransformer model still outperforms it by 13% in $R^{2}$ on the test set.
Our method significantly outperforms the deep hybrid models relying solely on single-region representations across both prediction tasks, demonstrating the substantial accuracy gains achieved by leveraging global urban information. Notably, GeoTransformer’s retrieval mechanism, which captures urban structural information without relying on predefined urban structures, exhibits clear superiority over GAT models that utilize graph structures built from the same information. This highlights the effectiveness of GeoTransformer in dynamically modeling urban dependencies and extracting meaningful global context for enhanced predictive performance.
Qualitative Analysis
To investigate the functionality of the dependency retrieval module, we visualize several top retrieval results for the target region. Figure 4 presents an example of a satellite image of the target region in the northwest area of the Auburn Gresham community and six top relevant regions based on the sparse matrix. The relevant regions span two major urban area types. The first includes functional zones near the target, such as local middle schools, parks, and churches (Region a-d in Figure 4). The second type of area covers transportation hubs within the community, such as railway hub centers and main roads within the community (Region e and f in Figure 4). These results showcase our retrieval module’s ability to capture urban structures within the specific spatial scale using the sparse matrix.
In addition to balancing the inconsistency in the information effectiveness of individual representation samples during the encoding process, we also argue that higher information entropy represents more complex and diverse regions, which is demonstrated by comparing regional representations with high and low information entropy (figure 5).
Figure 4: An example of dependency retrieval results.
Figure 5: A comparison of high and low information entropy regions.
@@
To investigate the necessity and effectiveness of each design in GeoTransformer, we conduct ablation experiments for the dependency retrieval module and two weighting methods for the geospatial attention module. The experiment setup and results for GDP prediction are shown in Table 3.
Four models were designed for the ablation study to evaluate the effectiveness of each component in GeoTransformer. Model 1 uses randomly selected regions without any attention weighting, serving as a baseline to demonstrate the importance of retrieval and attention weighting mechanisms. Model 2 applies the dependency retrieval mechanism on Model 1 to select relevant regions. Model 3 builds upon Model 2 by introducing spatial proximity weighting. Finally, Model 4 represents the complete GeoTransformer design, integrating the dependency retrieval module and both spatial proximity weighting and information entropy weighting to refine the attention mechanism.
The experimental results, presented in Table 3, show a clear and consistent improvement in performance with each added component. Model 1, with random selection, yields an $R_{2}$ near zero, indicating a chaotic model, while Model 2 introduces spatial dependencies through the retrieval mechanism, achieving a dramatic leap to over 80% $R_{2}$, demonstrating the effectiveness of the retrieval mechanism in selecting relevant regions. The addition of spatial proximity weighting in Model 3 further enhances accuracy, validating the importance of incorporating spatial relationships into the attention module. Finally, Model 4 achieves the best performance across all metrics, confirming the necessity of applying information entropy weighting.
@@
GeoTransformer demonstrates robust capabilities in urban forecasting, establishing a novel paradigm for region-based representation methods without reliance on predefined urban structural information. The experimental results validate the capacity of GeoTransformer, with ablation studies further highlighting the effectiveness of its key components. GeoTransformer still has limitations as it focuses solely on static validation and does not explore time-series tasks. However, the combination of the novel retrieval mechanism and attention-based model lays a strong foundation for future advancements in spatiotemporal forecasting.